A sample of an ideal gas has a volume of 3.10 L at 14.00*C and 180 atm. What is the volume of the gas at 21.80*C and 0.991 atm?

V=___ L
Answer #1

According to combined gas laws,

P1V1/T1 = P2V2/T2

P1 & P2 = Pressure

V1 & V2 = volume

T1 & T2 = Temperature

On applying combine gas law,

P1 = 180 atm. V1 = 3.10 L. T1 = 14^C

P2 = 0.991 atm. V2 = ? T2 = 21.80°C

( 180×3.10 )/14 = ( 0.991 × V2 )/21.80

V2 = ( 180 × 3.10 × 21.80 )/ 14 × 0.991

V2 = 876.78 L

Volume = 876.78 L